OpenStudy (anonymous):

Choose the ordered pair that is the solution to the system of equations. -5y + 4x = 8 -15y + 3x = -3 (3,4/5) (-3,-4) (-3,4/3) (3,0)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

the rest dont make scene to be correct

OpenStudy (anonymous):

thankz

OpenStudy (anonymous):

remember to medal

OpenStudy (anonymous):

need anymore help?

OpenStudy (anonymous):

no that was my only struggle

OpenStudy (mishary):

-5y+x=-1 x=-1+5y -5y+4x=8 Sub for x -5y+4(-1+5y)=8 -5y-4+20y=8 15y=12 Y=12/15=4/5 => x = -1+5(4/5)=-1+4=3